OpenStudy (anonymous):

Solve for n: n - (-8) = 12

OpenStudy (anonymous):

n - (-8) = n+8 so n+ 8 = 12 Subtract 8 from both sides, you get n = 4.

OpenStudy (anonymous):

Couldnt do it better my self nice job

OpenStudy (anonymous):

thank you!